Find the area of the region lying above X-axis and included between the circleΒ 

x2 + y2 = 2ax and the parabola y2 = ax lying above X-axis


1
Expert's answer
2021-04-29T12:03:30-0400

Solution.

"y^2+x^2=2ax,\newline
y^2=ax."

Let be "S" is finding area.

The required area at a = 1 (for a example):



Find the points of intersection:

"(x-a)^2+y^2-a^2=y^2-ax,"

"x^2-ax=0,"

"x(x-a)=0."

From here "(0,0)" and "(a,a)" are the points of intersection.


"S=\frac{1}{4}πr^2-\int_0^a \sqrt{ax}dx=\newline
=\frac{1}{4}πa^2-\frac{2}{3}a^2=\newline
=\frac{3π-8}{12}a^2."

Answer. "S=\frac{3π-8}{12}a^2."